OpenStudy (meilendurcer):

-3 1/5 divide 2/15= A. 1/24 B. -1/24 C. 24 D. -24

OpenStudy (anonymous):

first convert -3 1/5 into an improper fraction.

OpenStudy (meilendurcer):

how?

OpenStudy (anonymous):

-3 1/5 is the same thing as\[-3-\frac{1}{5}\]can you do this subtraction?

OpenStudy (meilendurcer):

8

OpenStudy (anonymous):

+8?

OpenStudy (meilendurcer):

no just 8

OpenStudy (anonymous):

8 and +8 are the same thing. not sure how you got that. Here's the first step:\[-3-\frac{1}{5}=\frac{-15}{5}-\frac{1}{5}\]

OpenStudy (meilendurcer):

ok now what? i got -8 was i suppose to?

OpenStudy (anonymous):

\[\frac{-15}{5}-\frac{1}{5}=\frac{-15-1}{5}=\frac{-16}{5}\]

OpenStudy (anonymous):

Next step is to do the division. \[\frac{-16}{5}\div \frac{2}{15}=\frac{-16}{5}\times \frac{15}{2}=...\]

OpenStudy (meilendurcer):

-96?

OpenStudy (anonymous):

not quite

OpenStudy (meilendurcer):

do i simplify?

OpenStudy (anonymous):